OpenStudy (anonymous):

separate the x ans y values with comma x=-42-8y -18x=-4-8y please i need help...

OpenStudy (anonymous):

So it says that x=-42-8y And they give you the other equation -18x=-4-8y So substitute what its given for to be, in this case x = -42 - 8y ..... into the equation -18x=-4-8y -18(-42-8y) = -4 - 8y